OpenStudy (anonymous):

If you were to use the substitution method to solve the following system, choose the new system of equations that would result if x was isolated in the first equation. x – 6y + 3z = 0 2x – 5y + 8z = 7 x + 7y – 4z = 13 –17y + 14z = 7 y – z = 13 7y + 2z = 7 13y – 7z = 13 13y + 7z = –13 –19y + 16z = –19 y – z = –13 9y + 0z = –19

OpenStudy (texaschic101):

x - 6y + 3z = 0 x = 6y - 3z 2x - 5y + 8z = 7 2(6y - 3z) - 5y + 8z = 7 12y - 6z - 5y + 8z = 7 7y + 2z = 7 x + 7y - 4z = 13 6y - 3z + 7y - 4z = 13 13y - 7z = 13 third one is your answer

OpenStudy (anonymous):

how did you figure it out because I tried solving and I got 7y+2z=7 -19+16z=-19 but that isn't an answer choice

OpenStudy (texaschic101):

well ...you got the first one For the second one....I just subbed 6y - 3z in for x in the third equation.

OpenStudy (anonymous):

where did u get it from?

OpenStudy (texaschic101):

get what from ?

OpenStudy (texaschic101):

I don't understand what you are asking

OpenStudy (texaschic101):

did you get 6y - 3z for x ?

OpenStudy (anonymous):

no?

OpenStudy (texaschic101):

you have to isolate x from the first equation x - 6y + 3z = 0 first add 6y to both sides x - 6y + 6y + 3z = 0 + 6y combine like terms x + 3z = 6y then you will subtract 3z from both sides to get x by itself x + 3z - 3z = 6y - 3z combine like terms x = 6y - 3z do you understand ?

OpenStudy (anonymous):

ok yeah

OpenStudy (texaschic101):

once you get x by itself, then you sub in what you got for x in the second equation, then in the third equation...does that make sense to you ?

OpenStudy (anonymous):

yes thanks

OpenStudy (texaschic101):

your very welcome .....any more questions ?

OpenStudy (anonymous):

are you from texas?

OpenStudy (texaschic101):

yes

OpenStudy (anonymous):

ok im done :) thanks

OpenStudy (texaschic101):

lol....no problem :)